A stone is fired with a velocity 100m/s in a direction making an angle 60°with the vertical. Calculate its maximum height attained and time of flight ?

Grade:11

2 Answers

Nikita arora
6 years ago
  • max height attained =v2sin2​​​angle /2g                                                                                                                                                     =   100×100×sin260 /2×10     =10000×3 /4 ×20    =375m                                                                                Time of flight =2vsinangle /g                                                                                                                                             = 2×100 ×sin 60 /10        =17.3 second
saniakhan
6 years ago
  • make data
    max height attained =v2sin2​​​angle /2g   
    Required
    Time..??
    Solution                                                                                                                                                  =   100×100×sin260 /2×10     =10000×3 /4 ×20    =375m                                                                                Time of flight =2vsinangle /g                                                                                                                                             = 2×100 ×sin 60 /10        =17.3 second
    Result=17.3 Seconds